Circular No. 7478

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000de IN NGC 4384 A. Dimai, Cortina d’Ampezzo, Italy, reports the discovery by M. Migli-

ardi, Tourtour, Haut-Provence, France, of an apparent supernova on CCD images taken on Aug. 10.9 and Aug. 11.9 UT (mag ∼ 16) with his 0.2-m f/10 Schmidt-Cassegrain telescope + CCD Hi-Sis 33 in the course of the CROSS program (cf. IAUC 7373). Confirmation was obtained on an image taken on Aug. 13.85 (limiting mag 18.5) with the Col Druscié Observatory 0.5-m Ullrich telescope that showed the supernova at mag ∼ 16. The new object is located at α = 12^{h}25^{m}12^{s}.17, δ = +54^{o}30′28′′.7 (equinox 2000), which is 1′′ east and 5′′ north of the nucleus of NGC 4384. A CROSS image of the same field on June 6 showed nothing at the position of SN 2000de (limiting mag ∼ 18.5). A. Nakamura, Kuma Kogen, Japan, forwards the following precise position end figures, measured by M. Aoki, Toyama, from five unfiltered CCD images taken with a 0.43-m f/6 reflector on Aug. 17.4, when the supernova was at mag 16.0: 12^{s}.20, 29′′.5. Nothing is visible on previous CCD images taken by Aoki between 1996 and 2000.

SUPERNOVA 2000df IN CGCG 51-70 M. Schwartz, Cottage Grove, OR, reports his discovery of an apparent

supernova (mag ∼ 17.2) on unfiltered CCD frames (limiting mag ∼ 18.5) taken on Aug. 17.204 and 18.194 UT, utilizing the Tenagra-I 0.35-m automated supernova patrol telescope (cf. IAUC 7456). The new object is located at α = 16^{h}14^{m}12^{s}.22, δ = +6^{o}09′00′′.6 (equinox 2000.0), which is ∼ 9′′.2 east and 2′′.6 south of the nucleus of CGCG 51-70. Nothing appears at the position of SN 2000df on red and blue plates of either Palomar Sky Survey.

COMET C/1999 T1 (McNAUGHT-HARTLEY) Visual m_{1} estimates: July 3.84 UT, 12.7 (M. Mattiazzo, Wallaroo, S.

Australia, 0.20-m reflector); 4.90, 12.6 (A. Pearce, Nedlands, W. Australia, 0.41-m reflector); 8.89, 12.3 (Pearce); Aug. 1.81, 10.9 (Mattiazzo).

COMETS DISCOVERED BY SOHO At the recent meetings of the IAU General Assembly it was decided

that comets discovered by SOHO could in the future be announced on the Minor Planet Electronic Circulars, a summary list then appearing on the IAU Circulars at a later date.

2000 August 18 \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT Elizabeth O. Waagen